Im very sorry i made a error on my last question the real question is how do you get k in this equation 6(k+2)+2(2k-5)=22
Answer by oscargut(682)

6(k+2)+2(2k-5)=22
then
6k+12+4k-10=22
then
10k+2=22
10k=20
k=2